OpenStudy (anonymous):

combine like terms. (-3y2 - 7y - 9) - (4y2 + 6y + 9)

OpenStudy (anonymous):

-3y^2 - 4y^2 = -7y^2 -7y + 6y = -y -9 -9 = -18 -7y^2 - y - 18 is your answer